Operations Report Apr 21st

Crew 245 Operations Report 21APR2021

SOL: 11

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7

From what time last night: 9:30 pm, SOC 60%

To what time this morning: 4:30 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 65% at 7:48 pm (Before generator is run at night)

Notes on power system: For the last several sols, the freezers in the Science Dome have been running in order to preserve science samples. Due to this, the station power consumption has been significantly higher than it previously was, resulting in the generator being turned on earlier (as early as 7:00 pm on colder days, and 9:00 pm on warmer days) once the SOC reaches below 70%.

Diesel Reading – 50%

Station Propane Reading – 70%

Water (loft tank): 25 gallons

Water Meter: 150665.5 units

Water (static tank): 135 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: No

Hours: 204.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Left tail-light on Perseverance rover is not functioning – noticed on EVA.

ATV’s Used: 350.2

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 0.5 gallon (estimated)

# Hours the ATVs were used today: 4 (but 0.5 hour of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew is rationing internet usage during the day to ensure that there is sufficient bandwidth for Mission Support Comms window.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 7, 9, and 11 used on EVA today, and airflow reduced for all suits approximately 3 hours into EVA. No further EVAs planned with these suits, but Crew Engineer suspects issue is with the chosen charging outlet – all three suits were charging from same power strip. Crew Engineer will test these suits tosol to determine root cause of issue and any potential fixes.

Summary of Hab operations: Flapping side panel on outside of Hab. Station Director secured panel with 2×4, so it is no longer making noise

Summary of GreenHab operations: Being used to host Crew Botanist regolith feasibility experiment, but no active operations.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

EVA Report – Apr 21st

Crew 245 EVA Report 21APR2021

EVA # 10

Author: Crew Scientist Olivia Ettlin

Purpose of EVA: Gathering soil samples and scouting for emergency shelters.

Start time: 1015 MDT

End time: 1330 MDT

Narrative: The EVA this morning started slightly after the beginning of the approved EVA window, as the crew was carefully gathering and sterilizing required scientific instruments according to XO Hume’s Planetary Protection Protocols. After we suited up and completed our 5-minute decompression in the airlock, we loaded our supplies into the Perseverance rover. Commander Dickstein and Health and Safety Officer Coultrup drove the Rover and Crew Scientist Ettlin took the ATV Serial Number 2. Starting charge on the Rover was 100% with 204.1 hours and the ATV required some choke before starting.

We traveled from the Hab to Entrance Road and proceeded south on Cow Dung Road for approximately 10 minutes until we reached Zubrin’s Head. Upon arrival at Zubrin’s head, we parked our vehicles on the side of the road. Using the analog map, we confirmed that our travel calculations had been accurate. From there, we embarked southwest on foot to explore the immediate region and investigate points of interest, which were communicated to us by our crewmates after their EVA to this site. As always, we carried the EVA med kit with us.

When we had completed gathering the samples, we returned to our vehicles and proceeded north on Cow Dung Road to Robert’s Rock Garden, where we did a final visit to drop waypoints and collect rock samples. From there, we returned to our vehicles and proceeded to Entrance Road, where we paused for approximately 20 minutes to capture some memorable photos with the MDRS entrance sign and facility as a backdrop, as directed by Mars Society Media Director Michael Stoltz. Finally, we returned back to the Hab. Upon returning, the Rover displayed 204.6 hours and 100% charge.

Destination: Zubrin’s Head

Coordinates of Destination: 12S 519500 UTM 4248100

Participants: Crew Scientist Olivia Ettlin, Commander Dylan Dickstein, and Health and Safety Officer Alex Coultrup

Road(s) and routes per MDRS Map: Hab to Entrance Rd, then South on Cow Dung road to Zubrin’s Head. Cow Dung Road to Entrance Rd, then back to the Hab.

Mode of travel: 2 team members in Perseverance Rover, 1 team member following close behind on ATV serial #2. All three members of the team walked the destination from the location where we parked the rovers safely on the side of the road.

Sol Summary Report Apr 21st

Crew 245 Sol Summary Report 21APR2021

Sol: 11

Summary Title: Approaching the Finish Line

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 08:30-09:30: Misc. breakfast and briefing

· 09:30-10:00: Prepare for EVA.

· 10:00-14:00: EVA.

· 14:00-15:00: Lunch! Jambalaya.

· 15:00-18:00: Separate science activities & rest time.

· 18:00-18:30: Crew Photos.

· 18:30-20:00: Dinner.

Look Ahead Plan: Tomorrow we have the final EVA of the mission – Crew #2.

Anomalies in work: None.

Weather: Nominal.

Crew Physical Status: No issues.

EVA: EVA #10 report coming in.

Reports to be filed: Operations Report, EVA Request, EVA Report

Support Requested: n/a

Copyright © The Mars Society. All rights reserved. | Main Site